# Locking-Free Training of Physics-Informed Neural Network for Solving Nearly Incompressible Elasticity Equations

### v1

@@ Abstract
Due to divergence instability, the accuracy of low-order conforming finite element methods for nearly incompressible homogeneous elasticity equations deteriorates as the Lamé coefficient $\lambda\to\infty$, or equivalently as the Poisson ratio $\nu\to 1/2$. This phenomenon, known as locking or non-robustness, remains not fully understood despite extensive investigation. In this paper, we propose a robust method based on a fundamentally different, machine-learning-driven approach. Leveraging recently developed Physics-Informed Neural Networks (PINNs), we address the numerical solution of linear elasticity equations governing nearly incompressible materials. The core idea of our method is to appropriately decompose the given equations to alleviate the extreme imbalance in the coefficients, while simultaneously solving both the forward and inverse problems to recover the solutions of the decomposed systems as well as the associated external conditions. Through various numerical experiments, including constant, variable and parametric Lamé coefficients, we illustrate the efficiency of the proposed methodology.
@@ 1 Introduction
$-\nabla\cdot\boldsymbol{\sigma}\bigl{(}{\bf u}({\boldsymbol{x}})\bigr{)}\,=\,{\bf f}({\boldsymbol{x}}),\quad{\boldsymbol{x}}\in\Omega,$ (1.1)
@@
[…]
$0<\mu_{\min}\leq\mu({\boldsymbol{x}})\leq\mu_{\max}<\infty~{}~{}{\rm and}~{}~{}0<\lambda_{\min}\leq\lambda({\boldsymbol{x}})\leq\lambda_{\max}<\infty,\quad{\boldsymbol{x}}\in\Omega.$
@@
For constant $\lambda$ and $\mu$, volumetric locking in standard conforming FEMs for the nearly incompressible elasticity problem (1.1) has catalyzed the formulation of numerous locking-free schemes. Nonconforming and mixed conforming-nonconforming FEMs (e.g., Arnold et al. [3]; Brenner and Sung [8]; Falk [17]; Gopalakrishnan and Guzmán [18]; Lee, Lee, and Sheen [27]; Mao and Chen [29]), weak Galerkin approaches (e.g., Chen and Xie [11]; Huo et al. [22]; Liu and Wang [28]), discontinuous Galerkin formulations (e.g., Bramwell et al. [7]; Di Pietro and Nicaise [14]; Hansbo and Larson [21]; Cockburn et al. [37]), and virtual element methods (e.g., Beirão da Veiga et al. [13]; Edoardo et al. [15]; Zhang et al. [39]) have successfully averted locking, albeit with greater algorithmic and analytical complexity. Extending many of these techniques to heterogeneous materials, where the Lamé coefficients vary with ${\boldsymbol{x}}$, introduces additional challenges. In [30], the authors address this by substituting the Lamé parameter $\lambda$ in the stiffness assembly with a mesh-dependent surrogate $\lambda_{h}=\frac{\lambda\,\mu L}{\mu L+\lambda\,h},$ where $h$ denotes the maximum finite element mesh size and $L$ is the diameter of $\Omega$. This simple yet effective modification preserves the efficiency of the lowest-order conforming FEMs while substantially alleviating volumetric locking.
@@
Deep artificial neural networks (DNNs) have garnered substantial attention as solvers for both forward and inverse partial differential equations (PDEs), owing to their capacity for high-dimensional approximation, the integration of automatic differentiation, and the maturity of open-source machine learning infrastructures. In particular, Physics-Informed Neural Networks (PINNs) incorporate the governing PDE residual directly into the loss functional, yielding a mesh-free paradigm that has experienced exponential growth in applications, surpassing 1 300 publications in 2021 alone [12]. Comprehensive reviews have critically examined the theoretical foundations, algorithmic limitations, and emerging use cases of PINNs [6, 9, 24], while advancing variants such as the variational hp-VPINN framework [25], the variable-scaling PINN [26], and physics-constrained neural formulations [38, 40] have demonstrated enhanced approximation fidelity and stability in challenging regimes.
@@
PINNs were introduced in $2017$ by Raissi et al. as a new class of data-driven solvers in a two-part article [32, 33]. Later on, and for solving nonlinear PDEs, the same authors created PINNs [34] which can handle both forward problems of estimating the solutions of governing mathematical models and inverse problems. PINNs approximate solutions by training a neural network to minimize a loss function; it includes terms reflecting both initial and boundary conditions in addition to the PDE residual at selected points in the domain (so-called collocation points). Incorporating a network for the PDE residual that encodes the governing physics equations is a significant novelty with PINNs. For a given input in the integration domain, PINNs produce an estimated solution of a differential equation after training. The basic concept behind PINN training is that it can be thought of as an unsupervised strategy that does not require labelled data, such as results from prior simulations or experiments. The PINN algorithm is essentially a mesh-free technique that finds PDE solutions by converting the problem of directly solving the governing equations into a loss function optimization problem. Owing to this, PINNs can handle PDEs in domains with complicated geometries or in very high dimensions. PINNs allow solutions to be made differentiable using analytical gradients and provide an easy way to solve forward and inverse problems using the same optimization problem (with minimal modifications). These can be considered some of the advantages of PINNs over conventional methods.
In the past few years, PINNs were applied by various authors to solve two-dimensional elasticity equations with constant Lamé coefficients, focusing on the case where the material is compressible, see [1, 10, 16, 19, 20, 35]. However, given the high representational capacity of neural networks used as the solution ansatz, PINNs suffer from locking (non-robustness) when applied to elasticity problems that are nearly incompressible, similar to what occurs in some traditional numerical methods. In this regime, assuming constant Lamé coefficients $\lambda$ and $\mu$, the fact that $\lambda\gg\mu$ leads to a highly imbalanced loss function. This imbalance causes the training dynamics to be dominated by the volumetric constraint, making the optimization landscape extremely ill-conditioned. Specifically, the role of the $\mu$-term, which governs shear deformation, is neglected during the training process. As a result, the neural network struggles to reduce the loss meaningfully and fails to learn an accurate solution, exhibiting locking behavior. In other words, the PINN-based locking originates from optimization issues arising from extreme parameter scaling. For constant $\lambda$ and $\mu,$ this phenomenon is illustrated in Figure 1, where we train neural networks to solve the linear elasticity equations using PINNs with different values of $\lambda$ and with $\mu=1.$
@@
The aim of the present work is to explore the application of PINNs for efficiently solving the elasticity equation (1.1)-(1.2) for the case of nearly incompressible materials in two and three dimensions, including constant (homogeneous material), variable (inhomogeneous material), and random Lamé coefficients. To overcome the non-robustness phenomena of PINNs in this case, we decompose the elasticity equation (1.1) into four different coupled PDEs. The key idea behind the decomposition is based on separating the strain tensor ${\boldsymbol{\varepsilon}}$ and the divergence $(\nabla\cdot){\bf I}$ operators. Then, the solution of the elasticity equation can be obtained via PINNs by minimizing the network’s loss function that comprises the residual error of the four governing PDEs and the boundary conditions. A detailed description of the proposed method, together with a brief summary of the PINN framework, will be discussed in the following two sections. Section 4 is devoted to illustrating the good performance of the proposed modified PINNs on various sample test problems of the form (1.1)-(1.2) in both two and three dimensions. Moreover, we will conduct the experiment for the case of variable Lamé, as well as the case of uncertainties in the Lamé coefficients. Finally, some concluding remarks will follow at the end of the paper.
@@ 2 Physics-informed neural networks
In this section, we provide a concise overview of the PINN framework, along with its extensions for parametric problems, which will be employed later in the paper. We begin by introducing the standard formulation of PINNs, which serves as the baseline approach [34]. Then, we introduce the parametric PINN, a novel approach that facilitates rapid inference of solutions under problems with varying coefficients, such as variable PDE coefficients, boundary conditions, initial conditions, or external forcing, without necessitating retraining. In this work, the Young’s modulus $E$ and Poisson’s ratio $\nu$ will be used as input coefficients, which determine the given elasticity equations. A comparative illustration of these methodologies is presented in Figure 2.
@@ 2.1 Vanilla PINN
First, we shall provide a brief description of PINNs [34]. Let us consider the following problem defined on a bounded open domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 1$):
@@
where $\mathcal{D}$ is a given differential operator and ${\bf u}:\overline{\Omega}\rightarrow\mathbb{R}^{m}$ (for some $m\in\mathbb{N}$) is the unknown solution, which is possibly multi-dimensional. Indeed, for time-dependent problems, the formulation of PINNs can be simplified by treating the temporal variable $t$ as an additional coordinate of ${\boldsymbol{x}}$, that is, $\Omega$ may represent the spatio-temporal domain. The core concept of PINNs is to approximate the solution ${\bf u}({\boldsymbol{x}})$ using a neural network ${\bf u}({\boldsymbol{x}};{\boldsymbol{\theta}})$, where ${\boldsymbol{\theta}}$ denotes the set of all coefficients. The model is trained by minimizing a physics-informed loss function, which includes the residual of the given PDE within the domain and the discrepancies between the neural network and the boundary conditions. More specifically, the neural network is trained to minimize the total loss
$\mathcal{L}_{\rm total}({\boldsymbol{\theta}})=\delta_{r}\mathcal{L}_{\rm res}({\boldsymbol{\theta}})+\delta_{b}\mathcal{L}_{\rm data}({\boldsymbol{\theta}}),$ (2.2)
@@
$\mathcal{L}_{\rm res}({\boldsymbol{\theta}})=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{D}[{\bf u}](\boldsymbol{x}_{r}^{i};{\boldsymbol{\theta}})-{\bf f}(\boldsymbol{x}_{r}^{i})\right|^{2},\quad\mathcal{L}_{\rm data}({\boldsymbol{\theta}})=\frac{1}{N_{b}}\sum_{j=1}^{N_{b}}\left|{\bf u}(\boldsymbol{x}_{b}^{j};{\boldsymbol{\theta}})-{\bf{g}}(\boldsymbol{x}_{b}^{j})\right|^{2}.$ (2.3)
@@ 2.2 Parametric PINN
Consider a parameter space $\mathcal{P}\subset\mathbb{R}^{k}$ for some $k\in\mathbb{N}$, and let $\boldsymbol{p}\in\mathcal{P}$ denote a vector of coefficients that governs the characteristics of the associated PDE problem. In particular, in the later numerical experiments, we will design a model that utilizes the coefficients, including Young’s modulus and Poisson’s ratio which determine the Lamé coefficients of the governing equations, as input features. This allows the model to rapidly predict solutions without retraining, even when these coefficients vary. The principal novelty behind the parametric PINN framework is the incorporation of problem-defining coefficients as additional inputs to the neural network. This extension allows the training process to be conducted over both the spatial domain $\Omega$ and the parameter space $\mathcal{P}\subset\mathbb{R}^{k}$, enabling the network to infer solutions that vary with respect to both spatial variables and coefficients.
To elaborate further, consider a parameter vector $\boldsymbol{p}\in\mathcal{P}$ which defines the PDE data, such as initial conditions, boundary conditions and source terms. In standard PINNs, one seeks to minimize the physics-informed loss function
$\min_{\boldsymbol{\theta}}\mathcal{L}_{\mathrm{total}}(\boldsymbol{\theta};\boldsymbol{p}),$
@@
$\mathcal{L}_{\mathrm{para}}(\boldsymbol{\theta}):=\int_{\mathcal{P}}\mathcal{L}_{\mathrm{total}}(\boldsymbol{\theta};\boldsymbol{p})\,\mathrm{d}\eta(\boldsymbol{p}),$

### v2

# A decomposition-based robust training of physics-informed neural networks for nearly incompressible linear elasticity

@@ Abstract
Due to divergence instability, the accuracy of low-order conforming finite element methods for nearly incompressible elasticity equations deteriorates as the Lamé coefficient $\lambda\to\infty$, or equivalently as the Poisson ratio $\nu\to 1/2$. This phenomenon, known as locking or non-robustness, remains not fully understood despite extensive investigation. In this work, we illustrate first that an analogous instability arises when applying the popular Physics-Informed Neural Networks (PINNs) to nearly incompressible elasticity problems, leading to significant loss of accuracy and convergence difficulties. Then, to overcome this challenge, we propose a robust decomposition-based PINN framework that reformulates the elasticity equations into balanced subsystems, thereby eliminating the ill-conditioning that causes locking. Our approach simultaneously solves the forward and inverse problems to recover both the decomposed field variables and the associated external conditions. We will also perform a convergence analysis to further enhance the reliability of the proposed approach. Moreover, through various numerical experiments, including constant, variable and parametric Lamé coefficients, we illustrate the efficiency of the proposed methodology.
@@ 1 Introduction
$\nabla\cdot\boldsymbol{\sigma}\bigl({\bf u}({\boldsymbol{x}})\bigr)\,=\,{\bf f}({\boldsymbol{x}}),\quad{\boldsymbol{x}}\in\Omega,$ (1.1)
@@
In this formulation, $\lambda$ and $\mu$ are the Lamé coefficients, which can be constants, functions of ${\boldsymbol{x}}\in\Omega$, or random fields, ${\bf f}:\Omega\to\mathbb{R}^{d}$ is the prescribed body force per unit volume, ${\bf{g}}:\Gamma\to\mathbb{R}^{d}$ is the given boundary condition, and ${\bf I}\in\mathbb{R}^{d\times d}$ is the identity tensor. The differential operators $\nabla$ (gradient) and $\nabla\cdot$ (divergence) act with respect to the spatial variable ${\boldsymbol{x}}\in\Omega$. To guarantee well-posedness of the problem (1.1)-(1.2), we assume that
$0<\mu_{\min}\leq\mu\leq\mu_{\max}<\infty~~{\rm and}~~0<\lambda_{\min}\leq\lambda\leq\lambda_{\max}<\infty\quad\text{on }\Omega.$
@@
For constant $\lambda$ and $\mu$, volumetric locking in standard conforming FEMs for the nearly incompressible elasticity problem (1.1) has catalyzed the formulation of numerous locking-free schemes. Nonconforming and mixed conforming-nonconforming FEMs (e.g., Arnold et al. [3]; Brenner and Sung [8]; Falk [17]; Gopalakrishnan and Guzmán [19]; Lee, Lee, and Sheen [28]; Mao and Chen [30]), weak Galerkin approaches (e.g., Chen and Xie [11]; Huo et al. [23]; Liu and Wang [29]), discontinuous Galerkin formulations (e.g., Bramwell et al. [7]; Di Pietro and Nicaise [14]; Hansbo and Larson [22]; Cockburn et al. [38]), and virtual element methods (e.g., Beirão da Veiga et al. [13]; Edoardo et al. [15]; Zhang et al. [40]) have successfully averted locking, albeit with greater algorithmic and analytical complexity. Extending many of these techniques to heterogeneous materials, where the Lamé coefficients vary with $\boldsymbol{x}$}, introduces additional challenges. In [31], the authors address this by substituting the Lamé parameter $\lambda$ in the stiffness assembly with an appropriate mesh-dependent surrogate $\lambda_{h}$ (with $\lambda_{h}<\lambda$), where $h$ denotes the maximum finite element mesh size and $L$ is the diameter of $\Omega$. This simple yet effective modification preserves the efficiency of the lowest-order conforming FEMs while substantially alleviating volumetric locking.
@@
Deep artificial neural networks (DNNs) have garnered substantial attention as solvers for both forward and inverse partial differential equations (PDEs), owing to their capacity for high-dimensional approximation, the integration of automatic differentiation, and the maturity of open-source machine learning infrastructures. In particular, Physics-Informed Neural Networks (PINNs) incorporate the governing PDE residual directly into the loss functional, yielding a mesh-free paradigm that has experienced exponential growth in applications, surpassing 1,300 publications in 2021 alone [12]. Comprehensive reviews have critically examined the theoretical foundations, algorithmic limitations, and emerging use cases of PINNs [6, 9, 25], while advancing variants such as the variational hp-VPINN framework [26], the variable-scaling PINN [27], and physics-constrained neural formulations [39, 41] have demonstrated enhanced approximation fidelity and stability in challenging regimes.
@@
PINNs were introduced in $2017$ by Raissi et al. as a new class of solvers in a two-part article [33, 34]. Later on, and for solving nonlinear PDEs, the same authors created PINNs [35], which can handle both forward problems of estimating the solutions of governing mathematical models and inverse problems. PINNs approximate solutions by training a neural network to minimize a loss function; it includes terms reflecting both initial and boundary conditions in addition to the PDE residual at selected points in the domain (so-called collocation points). Incorporating a network for the PDE residual that encodes the governing physics equations is a significant novelty with PINNs. For a given input in the integration domain, PINNs produce an estimated solution of a differential equation after training. The basic concept behind PINN training is that it can be thought of as an unsupervised strategy that does not require labelled data, such as results from prior simulations or experiments. The PINN algorithm is essentially a mesh-free technique that finds PDE solutions by converting the problem of directly solving the governing equations into a loss function optimization problem. Owing to this, PINNs can handle PDEs in domains with complicated geometries or in high dimensions. PINNs allow solutions to be made differentiable using analytical gradients and provide an easy way to solve forward and inverse problems using the same optimization problem (with minimal modifications). These can be considered some of the advantages of PINNs over conventional methods.
Despite the growing success of PINNs in solving a wide range of PDEs, including two-dimensional elasticity equations with constant Lamé coefficients [1, 10, 16, 20, 21, 36], their application to nearly incompressible elasticity problems remains highly challenging. Similar to classical low-order conforming FEMs, PINNs suffer from locking or divergence instability. This instability manifests as a severe loss of accuracy and robustness, preventing the network from properly learning the displacement and stress fields. Our investigations reveal that the root cause of this issue lies in the intrinsic imbalance of the governing equations, where the volumetric ($\lambda\,\bigl(\nabla\cdot{\bf u}\bigr)\,{\bf I}$) and deviatoric ($2\mu\,{\boldsymbol{\varepsilon}}({\bf u})$) terms of the stress tensor are scaled by vastly different magnitudes as the ratio of the Lamé coefficients $\lambda/\mu$ becomes very large. For constant $\lambda$ and $\mu,$ this phenomenon is illustrated in Figure 1, where we train neural networks to solve the linear elasticity equations using PINNs with different values of $\lambda$ and with $\mu=1.$ A more detailed discussion of this phenomenon will be provided in Section 3.
@@
In this work, and to overcome the aforementioned challenge in PINNs, we introduce a novel decomposition-based PINN framework that reformulates the nearly incompressible elasticity equations into balanced coupled subsystems. This decomposition effectively mitigates the ill-conditioning that causes locking and allows the network to learn each component with comparable accuracy. The solution can be obtained via PINNs by minimizing the network’s loss function that comprises the residual error of the decomposed governing PDEs and the boundary conditions. The proposed approach provides a stable and efficient alternative to standard PINNs, capable of accurately capturing the behavior of nearly incompressible materials in two and three dimensions, including constant (homogeneous material), variable (inhomogeneous material), and random Lamé coefficients.
The remainder of this paper is organized as follows. Section 2 provides a brief overview of the PINN frameworks, including its parametric extension, which serves as the foundation of our approach. In Section 3, we present a heuristic explanation of why the standard PINN fails to solve linear elasticity when $\lambda$ is large. Section 4 then describes the proposed method to resolve this problem in detail, followed by Section 5, which establishes a theoretical convergence analysis. Section 6 is devoted to demonstrating the effectiveness of the proposed modified PINNs through a series of numerical experiments on representative problems of the form (1.1)–(1.2) in both two and three dimensions. Moreover, we will conduct the experiment for the case of variable Lamé, as well as the case of uncertainties in the Lamé coefficients. Finally, some concluding remarks will follow in Section 7.
@@ 2 Physics-informed neural networks
In this section, we provide a concise overview of the PINN framework, along with its extensions for parametric problems, which will be employed later in the paper. We begin by introducing the standard formulation of PINNs, which serves as the baseline approach [35]. Then, we introduce the parametric PINN, a novel approach that facilitates rapid inference of solutions under problems with varying parameters, such as variable PDE coefficients, boundary conditions, initial conditions, or external forcing, without necessitating retraining. In this work, the Young’s modulus $E$ and Poisson’s ratio $\nu$ will be used as input coefficients, which determine the given elasticity equations. A comparative illustration of these methodologies is presented in Figure 2.
@@ 2.1 Vanilla PINN
First, we shall provide a brief description of PINNs [35]. Let us consider the following problem defined on a bounded open domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 1$) with a boundary $\Gamma=\partial\Omega$:
@@
where $\mathcal{D}$ is a given differential operator and ${\bf u}:\overline{\Omega}\rightarrow\mathbb{R}^{m}$ (for some $m\in\mathbb{N}$) is the unknown solution, which is possibly multi-dimensional. Indeed, for time-dependent problems, the formulation of PINNs can be simplified by treating the temporal variable $t$ as an additional coordinate of $\boldsymbol{x}$}, that is, $\Omega$ may represent the spatio-temporal domain. The core concept of PINNs is to approximate the solution ${\bf u}({\boldsymbol{x}})$ with a neural network ${\bf u}({\boldsymbol{x}};{\boldsymbol{\theta}})$, where $\boldsymbol{\theta}$} denotes the trainable parameters which will be described in more detail in Section 5, and to learn $\boldsymbol{\theta}$} by minimizing a loss that enforces the governing PDE residual in the domain $\Omega$ together with the boundary conditions on $\partial\Omega$. The model is trained by minimizing a physics-informed loss function, which includes the residual of the given PDE within the domain and the discrepancies between the neural network and the boundary conditions. More specifically, the neural network is trained to minimize the total loss
$\mathcal{L}({\boldsymbol{\theta}})=\delta_{r}\mathcal{L}_{r}({\boldsymbol{\theta}})+\delta_{b}\mathcal{L}_{b}({\boldsymbol{\theta}}),$ (2.2)
@@
$\mathcal{L}_{r}({\boldsymbol{\theta}})=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{D}[{\bf u}](\boldsymbol{x}_{r}^{i};{\boldsymbol{\theta}})-{\bf f}(\boldsymbol{x}_{r}^{i})\right|^{2},\quad\mathcal{L}_{b}({\boldsymbol{\theta}})=\frac{1}{N_{b}}\sum_{j=1}^{N_{b}}\left|{\bf u}(\boldsymbol{x}_{b}^{j};{\boldsymbol{\theta}})-{\bf{g}}(\boldsymbol{x}_{b}^{j})\right|^{2}.$ (2.3)
@@ 2.2 Parametric PINN
Consider a parameter space $\mathcal{P}\subset\mathbb{R}^{k}$ for some $k\in\mathbb{N}$, and let $\boldsymbol{p}\in\mathcal{P}$ denote a vector of coefficients that governs the characteristics of the associated PDE problem. In particular, in the later numerical experiments, we will design a model that utilizes the coefficients, including Young’s modulus $E$ and Poisson’s ratio $\nu$, which determine the Lamé coefficients of the governing equations, as input features. The principal novelty behind the parametric PINN framework is the incorporation of problem-defining coefficients as additional inputs to the neural network. This extension allows the training process to be conducted over both the spatial domain $\Omega$ and the parameter space $\mathcal{P}\subset\mathbb{R}^{k}$, enabling the network to infer solutions that vary with respect to both spatial variables and coefficients.
To elaborate further, consider a parameter vector $\boldsymbol{p}\in\mathcal{P}$ which defines the PDE data, such as initial conditions, boundary conditions, and source terms. In standard PINNs, one seeks to minimize the physics-informed loss function
$\min_{{\boldsymbol{\theta}}}\mathcal{L}({\boldsymbol{\theta}};\boldsymbol{p}),$
@@
$\mathcal{L}_{\mathrm{para}}(\boldsymbol{\theta}):=\int_{\mathcal{P}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{p})\,\mathrm{d}\eta(\boldsymbol{p}),$
@@
3 Effects of incompressibility: Volumetric locking in PINN training
In this section, we introduce a heuristic perspective on why the training of a standard PINN fails when applied to nearly incompressible linear elasticity equations. To simplify the heuristics, we assume that the material is homogeneous (i.e., $\lambda$ and $\mu$ are constant) and ${\bf{g}}={\bf 0}$. We consider a straightforward implementation of PINNs for linear elasticity, where, in this case, the loss function is of the form
$\mathcal{L}({\boldsymbol{\theta}})=$ $\frac{\delta_{r}}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{D}({\bf u})({\boldsymbol{x}}^{i}_{r};{\boldsymbol{\theta}})+{\bf f}({\boldsymbol{x}}^{i}_{r})\right|^{2}+\frac{\delta_{b}}{N_{b}}\sum_{j=1}^{N_{b}}\left|{\bf u}({\boldsymbol{x}}^{j}_{b};{\boldsymbol{\theta}})\right|^{2},$ (3.1)
where $\mathcal{D}({\bf u})=\nabla\cdot(\lambda\nabla\cdot{\bf u}{\bf I})+\nabla\cdot(\mu{\boldsymbol{\varepsilon}}({\bf u}))$.
Note that if we replace $\bf f$} in (1.1) by another function ${\bf f}_{2}$ which satisfies ${\bf f}_{2}({\boldsymbol{x}}_{r}^{i})={\bf f}({\boldsymbol{x}}_{r}^{i})$ we obtain the same loss function (3.1). However, the exact solution $\bf u$} of (1.1) with right-hand sides $\bf f$} differs from the exact solution ${\bf u}_{2}$ of (1.1) with right-hand side ${\bf f}_{2}$ in general. Hence, there is no unique solution $\bf u$} of (1.1) associated with the minimizer of the loss function (3.1). Hence, the discrete loss function (3.1) does not have a unique minimizer in general (for instance, both $\bf u$} and ${\bf u}_{2}$ minimize the loss function). In the following, we illustrate that, when $\lambda$ is large, the set of minimizers of (3.1) contains a ‘solution’ which is close to zero, and that standard PINN approximates this solution.
When training a neural network using a gradient descent method, the starting point of the gradient descent method determines which minimizer the optimization procedure will converge to. We discuss this in the context where $\lambda$ becomes large. The gradient flow equation, given by
$\frac{\mathrm{d}}{\mathrm{d}t}{\boldsymbol{\theta}}(t)=-\nabla_{{\boldsymbol{\theta}}}\mathcal{L}({\boldsymbol{\theta}}),$
models the training behavior of the coefficients $\boldsymbol{\theta}$} of a neural network when the step-size of the training goes to 0 such that the parameters $\boldsymbol{\theta}$} depend continuously on the training time $t$. Explicitly,
$\nabla_{{\boldsymbol{\theta}}}\mathcal{L}({\boldsymbol{\theta}})$ $=\frac{2\delta_{r}}{N_{r}}\sum_{i=1}^{N_{r}}\left(\mathcal{D}({\bf u})({\boldsymbol{x}}^{i}_{r};{\boldsymbol{\theta}})+{\bf f}({\boldsymbol{x}}^{i}_{r})\right)\mathcal{D}(\nabla_{{\boldsymbol{\theta}}}{\bf u})({\boldsymbol{x}}_{r}^{i};{\boldsymbol{\theta}})+\frac{2\delta_{b}}{N_{b}}\sum_{j=1}^{N_{b}}{\bf u}({\boldsymbol{x}}^{j}_{b};{\boldsymbol{\theta}})\nabla_{{\boldsymbol{\theta}}}{\bf u}({\boldsymbol{x}}_{b}^{j};{\boldsymbol{\theta}})$ (3.2)
$=\lambda^{2}{\bf a}+\lambda\mu{\bf b}+\mu^{2}{\bf c}+{\bf d},$
where
${\bf a}$ $=\frac{2\delta_{r}}{N_{r}}\sum_{i=1}^{N_{r}}[\nabla\cdot(\nabla\cdot({\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}}){\bf I})][\nabla\cdot(\nabla\cdot(\nabla_{{\boldsymbol{\theta}}}{\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}})){\bf I})],$
${\bf b}$ $=\frac{2\delta_{r}}{N_{r}}\sum_{i=1}^{N_{r}}[\nabla\cdot(\nabla\cdot({\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}}){\bf I})][\nabla\cdot{\boldsymbol{\varepsilon}}(\nabla_{{\boldsymbol{\theta}}}{\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}}))]+[\nabla\cdot(\nabla\cdot(\nabla_{{\boldsymbol{\theta}}}{\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}})){\bf I})][\nabla\cdot{\boldsymbol{\varepsilon}}({\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}}))],$
${\bf c}$ $=\frac{2\delta_{r}}{N_{r}}\sum_{i=1}^{N_{r}}[\nabla\cdot{\boldsymbol{\varepsilon}}({\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}}))][\nabla\cdot{\boldsymbol{\varepsilon}}(\nabla_{{\boldsymbol{\theta}}}{\bf u}({\boldsymbol{x}}_{i}^{e};{\boldsymbol{\theta}}))],$
${\bf d}$ $=\frac{2\delta_{b}}{N_{b}}\sum_{j=1}^{N_{b}}{\bf u}({\boldsymbol{x}}^{j}_{b};{\boldsymbol{\theta}})\nabla_{{\boldsymbol{\theta}}}{\bf u}({\boldsymbol{x}}_{b}^{j};{\boldsymbol{\theta}}).$
Assume that the sampling points ${\boldsymbol{x}}_{r}^{i},{\boldsymbol{x}}_{b}^{j}$ and the initial neural network for the PINN training have been chosen randomly, irrespective of the values of $\lambda$ and $\mu$. [^1] Note that the solution $\bf u$$^{\ast}$ of the linear elasticity equation changes if we change $\lambda$ and $\mu$. This random choice induces distributions on the vectors ${\bf a},{\bf b},{\bf c},{\bf d}$, which are again independent of $\lambda$ and $\mu$. At the initialization of the gradient descent algorithm, the direction of the gradient of the loss function $\nabla_{{\boldsymbol{\theta}}}\mathcal{L}$ depends, in particular, on $\lambda,\mu,{\bf a},{\bf b},{\bf c},{\bf d}$. As $\lambda$ increases, the direction $\bf a$} becomes more dominant. To illustrate this, consider
$\frac{1}{\lambda^{2}}\nabla_{{\boldsymbol{\theta}}}\mathcal{L}={\bf a}+\frac{\mu}{\lambda}{\bf b}+\frac{\mu^{2}}{\lambda^{2}}{\bf c}+\frac{1}{\lambda^{2}}{\bf d}.$ (3.3)
As $\lambda$ increases, the direction of the gradient descent algorithm is increasingly aligned with the direction of $\bf a$} as the term $\frac{\mu}{\lambda}{\bf b}+\frac{\mu^{2}}{\lambda^{2}}{\bf c}$ becomes negligible (the importance of the boundary term $\frac{1}{\lambda^{2}}{\bf d}$ can be adjusted via the weight $\delta_{b}$). This means that with increasing probability, in the optimization procedure, $\boldsymbol{\theta}$} is modified in a direction which minimizes the loss function
$\overline{\mathcal{L}}=\frac{\delta_{r}}{N_{r}}\sum_{i=1}^{N_{r}}|\nabla\cdot(\nabla\cdot{\bf u}({\boldsymbol{x}}_{r}^{i};{\boldsymbol{\theta}}){\bf I})|^{2}+\frac{\delta_{b}}{N_{b}}\sum_{j=1}^{N_{b}}|{\bf u}({\boldsymbol{x}}_{b}^{j};{\boldsymbol{\theta}})|^{2},$ (3.4)
since $\nabla_{{\boldsymbol{\theta}}}\overline{\mathcal{L}}=\lambda^{2}{\bf a}+{\bf d}$. However, minimizing the loss function (3.4) corresponds to approximating a function $\bar{{\bf u}}$ which minimizes the loss function
$\delta_{r}\left\|\nabla\cdot(\nabla\cdot{\bf u}{\bf I})\right\|_{L_{2}(\Omega)}^{2}+\delta_{b}\left\|{\bf u}\right\|_{L_{2}(\partial\Omega}^{2}.$ (3.5)
However, it is easy to verify that the minimizer of (3.5) is the function which is $\bf 0$} everywhere. In other words, the basin of attraction for gradient descent methods for the $\bf 0$} solution increases with $\lambda$, so that for nearly incompressible materials the standard PINN leads to the $\bf 0$} solution unless the initial guess is already very close to the true solution. In particular, we observe this kind of behavior in the numerical experiments for large $\lambda$ and zero Dirichlet boundary conditions using a standard PINN approach in Section 6.1.
Figure 3: Cosine similarity between $\nabla_{{\boldsymbol{\theta}}}\mathcal{L}$ and $\nabla_{{\boldsymbol{\theta}}}\overline{\mathcal{L}}$ during training. This demonstrates that minimizing the loss $\mathcal{L}$ proceeds almost identically to minimizing the loss $\overline{\mathcal{L}}$, thereby explaining the occurrence of the locking phenomenon in PINN training.
To make this heuristic more concrete, we conducted a numerical experiment. Specifically, we initialized two neural networks identically and trained them separately using loss functions $\mathcal{L}$ in (3.1) and $\overline{\mathcal{L}}$ in (3.4) respectively, while observing their training dynamics. To examine the directions of the parameter updates during training, we recorded quantities $\nabla_{{\boldsymbol{\theta}}}\mathcal{L}$ and $\nabla_{{\boldsymbol{\theta}}}\overline{\mathcal{L}}$ at each epoch and measured the cosine similarity between them. Interestingly, as shown in Figure 3, the cosine similarity between $\nabla_{{\boldsymbol{\theta}}}\mathcal{L}$ and $\nabla_{{\boldsymbol{\theta}}}\overline{\mathcal{L}}$ remained close to 1 throughout the training process. This indicates that, when $\lambda$ is large, the training of a standard PINN proceeds almost as if it were minimizing the loss function $\overline{\mathcal{L}}$. However, since the solution obtained by minimizing $\overline{\mathcal{L}}$ differs from the true solution of the original PDE, this explains why the training of a standard PINN fails to converge properly in the nearly incompressible regime. In other words, when solving linear elasticity problems with large $\lambda$ using PINNs, an imbalance among the terms in the loss function leads to a phenomenon similar to locking, where the training does not proceed correctly. Much like in classical FEM, overcoming this locking effect remains a significant computational challenge. In the next section, we propose a method to address this issue, thereby enabling PINNs to effectively compute solutions in the nearly incompressible regime.